OpenStudy (anonymous):

simplify square root 50x^6

OpenStudy (lgbasallote):

is 50 in the square root as well?

OpenStudy (anonymous):

Yes:)

OpenStudy (lgbasallote):

sqrt (25 x 2 ) times sqrtx^6 5 sqrt 2 times x^6/2 5 sqrt 2 times x^3 so... 5x^3 sqrt 2

OpenStudy (anonymous):

sqrt(50 x^6) = sqrt(2 * 5^2 * (x^3)^2) = sqrt(2) * 5 * x^3 = 5 sqrt(2) x^3